OpenStudy (anonymous):

Given that f(x) = 3x + 1 and g(x) = the quantity of 4x plus 2 divided by 3 solve for g(f(0)).

OpenStudy (anonymous):

The best way to think about this problem is what are you plugging int to what? g(f(0)) means, you plug 0 into f(x). Whatever number you get from that, you plug into g(x)! Give it a try and let me know what you get

OpenStudy (anonymous):

so right now i got g(f(0)=\[\frac{ 4(3x+1)+2 }{ 3 }\]

OpenStudy (anonymous):

First let's find F(0). Plug in 0 for x in F(x) F(x) = 3x +1 F(0) = 3(0) + 1 F(0) = 0 +1 F(0) = 1 So now, you plug in 1 into g(x) G(1) = ?????

OpenStudy (anonymous):

G(x)=\[\frac{ 4x+2 }{ 3 }\] G(1)=\[\frac{ 4(1)+2 }{3}\] G(1)=\[\frac{ 6 }{ 3 }\] G(1)=\[\frac{ 2 }{ 1 }\] = 2 G(1)=2

OpenStudy (anonymous):

Exactly! So g(f(0)) = 2 Remember the procedure and you will do fine!

OpenStudy (anonymous):

g(f(0)) = g(1) = 2

OpenStudy (anonymous):

can you help me with this one Given f(x) = the quantity of 2x+4/ 5 solve for f-1(3).?

OpenStudy (anonymous):

I don't understand the very last part of your problem "f-1(3)" Can you rephrase that?

OpenStudy (anonymous):

idk that what the paper says i think it means the inverse